OpenStudy (anonymous):

solve the equation f(x)=x^2-5X-12 f(x)= f(-2)

OpenStudy (saifoo.khan):

f(x)=x^2-5X-12 f(x) = f(-2) we will write as x^2-5x-12 = x^2-5X-12 x^2-5x-12 = (-2)^2-5(-2)-12 And solve.

OpenStudy (anonymous):

thanks!